OpenStudy (anonymous):

can you help with this problem i deffiantly dont understand this 3x+12/4x-18 divided by x+4/x-3

OpenStudy (anonymous):

0.48

OpenStudy (anonymous):

@Kimmieboo1316 ! lol here you go!

OpenStudy (anonymous):

x=0.48

OpenStudy (mertsj):

Dividing is the same as multiplying by the reciprocal of the divisor. Factor everything that can be factored, cancel the common factors and multiply.

OpenStudy (anonymous):

thats not one of my answers :/

OpenStudy (mertsj):

\[\frac{3x+12}{4x-18} \times \frac{x-3}{x+4}\]

OpenStudy (anonymous):

@Mertsj this is @Kimmieboo1316 question help her if you can !

OpenStudy (anonymous):

thankd and yes thats the problem

OpenStudy (mertsj):

Pull a 3 out of 3x+12. Can you do that?

OpenStudy (anonymous):

im not sure how do to any of this

OpenStudy (mertsj):

3x+12=3(x+4)

OpenStudy (anonymous):

um okay?

OpenStudy (mertsj):